You’ve observed the following returns on SkyNet Data Corporation’s stock over the past five years: 17 percent, –15 percent, 19 percent, 29 percent, and 10 percent. Suppose the average inflation rate over this period was 2.6 percent and the average T-bill rate over the period was 4.3 percent.

a.

What was the average real return on the company's stock? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b.

What was the average nominal risk premium on the company's stock? (Do not round intermediate calculations and enter your answer as a percent rounded to 1 decimal place, e.g., 32.1.)
Answer #1

Average Nominal Return = [0.17 + (-0.15) + 0.19 + 0.29 + 0.10] / 5
Average Nominal Return = 0.60 / 5
Average Nominal Return = 0.12 or 12.00%

Answer a.

Average Real Return = (Average Nominal Return - Inflation Rate) / (1 + Inflation Rate)
Average Real Return = (0.12 - 0.026) / (1 + 0.026)
Average Real Return = 0.094 / 1.026
Average Real Return = 0.0916 or 9.16%

Answer b.

Average Nominal Risk Premium = Average Nominal Return - Average Risk-free Rate
Average Nominal Risk Premium = 12.00% - 4.30%
Average Nominal Risk Premium = 7.70%